A 12.0 kg object is pushed with a horizontal force of 6.0 N easy across a horizontal table. If the force of friction between the two surfaces is 2.0 N, what is the acceleration of the object?



1
Expert's answer
2021-05-18T11:10:21-0400
"F_{net}=ma,""F_h-F_{fr}=ma,""a=\dfrac{F_h-F_{fr}}{m}=\dfrac{6\ N-2\ N}{12\ kg}=0.33\ \dfrac{m}{s^2}."